OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4

hartnn (hartnn):

Equation of line with point x1,y1 and slope =m is \(y-y_1=m(x-x_1)\) just plug in values!

OpenStudy (anonymous):

Well to find the equation we use y - b = m(x-a) X and Y are your points (3, 2) and your slope is your M (gradient) :) So subbing in we get y - 2 = 4(x - 3) y - 2 = 4x - 12 y = 4x - 10 :) glad I could help.

OpenStudy (anonymous):

Thanks !